Sunny Side, Inc. granted its CEO 5,000 options on July 1, 2016 with an exercise price of $25. The market value of options based on an option pricing model is $300,000. The market price of the stock on July 1, 2016 was $30 per share. The vesting period is two years and options expire on July 1, 2021. The fiscal year end is December 31. If instead of issuing the options the company issued restricted stock with a three-year vesting period, then how many restricted shares would need to be issued to record the same amount of compensation expense for the year ended December 31, 2016.
Answer #1

Total amount of Compensation Expense = (30-25)*5000 Options =25000

Total vesting period =2 yrs

Restricted Shares total period = 3 yrs

Vesting Period from July 2016 to December 2016 = 6 months

Amount of Expense allocated to 6 months period is 25000/36months*6 months = 4166

No of restricted shares to be allotted=4166/30=139 Shares